OpenStudy (anonymous):

Write an equation for a line in slope-intercept form that passes through the points (3, 4) and (1, -2). y = x – 1 y = x + 1 y = 3x – 5 y = 3x + 5

OpenStudy (ash2326):

Equation of a line passing through two points (x1, y1) and (x2, y2) is given as: \[\frac{y-y2}{x-x2}=\frac{y2-y1}{x2-x1}\] here we have \[(x2, y2)=(3, 4)\] \[(x1, y1)=(1, -2)\] Now let's substitute these in the equation \[\frac{y-4}{x-3}=\frac{4-(-2)}{3-1}\] Just simplify and you'll get the equation of the line